# Periodic orbits for square and rectangular billiards

### v1

@@ 3.3 Singular orbits for the square billiard
Proof.
Suppose for the sake of contradiction that there exists a singular orbit that starts from a vertex with irrational slope. Without loss of generality, we may assume that this billiard trajectory starts at vertex A of the square ABCD; if not, we rotate the table and relabel the vertices. We now consider the corresponding unfolded trajectory in the plane $\mathbb{R}^{2}$. Since we assumed that the billiard trajectory is singular, the unfolded trajectory must pass through a shifted copy of one of the vertices; by Definition 3.1, these shifted copies lie in $\mathbb{Z}\times\mathbb{Z}$. Therefore, the unfolded trajectory contains a line segment from the original vertex A at $(0,0)$ to a shifted copy of this same or another vertex at the point, say, $(m,n)\in\mathbb{Z}\times\mathbb{Z}$; here both $n\neq 0$ and $m\neq 0$, because the slope is neither 00 nor $\infty$. This means that the unfolded trajectory is the straight line with slope $\frac{m}{n}\in\mathbb{Q}$. Any possible rotation back to the original billiard table either preserves this slope (rotation by 00 or $\pi$), or changes it to $\frac{n}{m}\in\mathbb{Q}$ (rotation by ±$\pm\frac{\pi}{2}$), both of which are rational; a contradiction. ∎
Corollary 3.8.
@@
Indeed, both billiard trajectories unfold to the same line with rational slope in the plane $\mathbb{R}^{2}$. Hence, every singular orbit is contained in a billiard trajectory that both starts and terminates at a vertex. Such billiard trajectories are called generalised diagonals (Katok 1987).
Definition 3.9.
@@
Proof.
The unfolded trajectory of such a generalised diagonal is the line through the origin with slope $\frac{m}{n}$ that passes through all vertex points $(i,j)\in\mathbb{Z}\times\mathbb{Z}$ that are integer multiples of $(n,m)$. There are no other points $(i,j)\in\mathbb{Z}\times\mathbb{Z}$ on this line, because $\gcd{(m,n)}=1$. From $(0,0)$ to $(n,m)$, we encounter exactly $m-1$ vertical reflections and exactly $n-1$ horizontal reflections, which means that there are $m-1$ collision points on the horizontal sides AB or CD and $n-1$ collision points on the vertical sides BC or DA. The length of this singular trajectory is, thus, $m-1+n-1=m+n-2$, as required. ∎
@@
Proof.
Since $\tan{(\alpha_{0})}=\frac{p}{q}$ is rational, this billiard trajectory cannot be a non-periodic orbit, because non-periodic orbits unfold to lines with irrational slopes; cf. Remark 2. If it is periodic, then it is of type $(p,q)$ and Definitions 2.2 and 3.4 imply that such a periodic orbit will be a member of the class $\mathcal{C}_{K}(p)$, with $K=p+q$. Furthermore, the billiard trajectory unfolds to a line (or line segment) with fixed slope $\frac{p}{q}$, so if it is singular, it must lie on a generalised diagonal with this slope. Since $\gcd{(p,q)}=2$, Proposition 3.10 implies that the generalised diagonal will have length $\frac{p}{2}+\frac{q}{2}-2$ as required, and it will, indeed, have $\frac{p}{2}-1$ collision points on the horizontal sides AB or CD and the remaining $\frac{q}{2}-1$ collision points on the vertical sides BC or DA. Hence, the proof is complete as soon as we show that a billiard trajectory generated by the pair $\langle P_{0},\alpha_{0}\rangle$ is singular if and only if $P_{0}=\frac{2\ell}{p}$ for $\ell=0,1,\ldots,\frac{p}{2}-1$.
@@ 5 Discussion
We completely classified the existence and nature of all periodic orbits for a rectangular billiard table with aspect ratio $1:\varrho$, where $0<\varrho<\infty$; we discussed the square billiard table in full detail, which is included in this setting as the case with $\varrho=1$. The class of rectangular billiars is special, because it is the only class of polygonal billards for which the classification of all periodic orbits is preserved under a non-uniform scaling (Calderon, Coles, Davis, Lanier and Oliveira 2018; Duchin, Erlandsson, Leininger and Sadanand 2019). The special property that the sides of a rectangular table are perpendicular to each other enables us to characterise the periodic orbits for this class of billard tables in unprecedented detail. In particular, we show that rectangular billiards admit period-$K$ orbits for any even period, but not for any odd period $K\in\mathbb{N}$. As soon as $K\geq 6$, there are different types of period-$K$ orbits, determined by the difference between the numbers of collisions with the two pairs of parallel sides. Each type is completely determined by the slope that the periodic orbit should have with respect to one of the sides of the rectangle. Furthermore, the periodic orbit can be realised with any initial point on this side, except for a finite set of exactly $\frac{1}{2}K-1$ points, other than the table corners, that give rise to singular orbits; the unfolding of such singular orbits leads to generalised diagonals in the plane. We defined equivalence classes for each type and proved that the total of different types of period-$K$ orbits, for any $K\in 2\,\mathbb{N}$ is given by Euler’s totient function $\varphi(N)$ evaluated at $N=\frac{1}{2}K$; the totients $n\in\{1,\ldots,N\}$ such that $\gcd{(n,N)}=1$ define pairs $(m,n)$ with $m=N-n$ that uniquely define the slopes needed to generate a periodic orbit of a specific type.
@@
As mentioned, the difference between the square and rectangular billiard is a stretching of one pair of parallel sides, which leads to an adjustment by the aspect-ratio parameter $\varrho$ of the slopes required to generate particular periodic orbits. It is tempting to apply a more general transformation, for example, one that scales and shears the sides, such that the billiard table becomes a parallelogram. Unfortunately, the unfolding technique generally fails to tile the plane for parallelogram billiards (Fox and Kershner 1936; Zemlyakov and Katok 1975; Tabachnikov 2005). We find that there is very little known regarding parallelogram billiards. Do parallelograms also only admit periodic orbits with even periods? What are necessary and sufficient conditions for existence of a period-$K$ orbit and how do we construct such a periodic orbit for the parallelogram?
@@
The non-existence of periodic orbits with odd periods has been proven for parallelograms consisting of two equilateral triangles glued together, that is, one angle is $\frac{\pi}{3}$ radians (Baer et al. 2019). Furthermore, it is known that periodic orbits for the parallelogram with angles $\frac{\pi}{4}$ and $\frac{3\pi}{4}$ do not persist under perturbations, that is, so-called stable periodic orbits do not exist for this parallelogram table (Rozikov 2018). More generally, it is known that polygonal billiard tables with angles equal to rational multiples of $\pi$ have many periodic trajectories (Boshernitzan et al. 1998; DeMarco 2011; Masur and Tabachnikov 2002; Tabachnikov 2005). However, almost nothing is known for billiard tables with irrational angles. For example, as we already alluded to in the introduction, it is an open question whether periodic orbits exist for triangular billiard tables with arbitrary (irrational) angles. The (rational) angles for acute triangles that lead to a lattice polygon, that is, triangles that unfold to a planar tiling, have all been listed (Kenyon and Smillie 2000; Puchta 2001). Hooper (2007) has shown that periodic orbits for right-triangle billiards are never stable. In contrast, if one perturbs the right triangle such that it has an obtuse angle, then it will always have a stable periodic orbit (Schwartz 2006, 2009); the proof is restricted to obtuse angles that do not exceed $\frac{5\pi}{9}$. For larger angles, we are only aware of existence results of periodic orbits for isosceles triangles (Hooper and Schwartz 2009). We believe that progress can be made from the study of periodic orbits in a parallelogram table, which we view as continuous deformations from corresponding periodic orbits in the rectangular table. Results in this direction are left for future work.

### v2

@@ 3.3 Singular orbits for the square billiard
Proof. Suppose for the sake of contradiction that there exists a singular orbit that starts from a vertex with irrational slope. Without loss of generality, we may assume that this billiard trajectory starts at vertex A of the square ABCD; if not, we rotate the table and relabel the vertices. We now consider the corresponding unfolded trajectory in the plane $\mathbb{R}^{2}$. Since we assumed that the billiard trajectory is singular, the unfolded trajectory must pass through a shifted copy of one of the vertices; by Definition 3.1, these shifted copies lie in $\mathbb{Z}\times\mathbb{Z}$. Therefore, the unfolded trajectory contains a line segment from the original vertex A at $(0,0)$ to a shifted copy of this same or another vertex at the point, say, $(m,n)\in\mathbb{Z}\times\mathbb{Z}$; here both $n\neq 0$ and $m\neq 0$, because the slope is neither 00 nor $\infty$. This means that the unfolded trajectory is the straight line with slope $\frac{m}{n}\in\mathbb{Q}$. Any possible rotation back to the original billiard table either preserves this slope (rotation by 00 or $\pi$), or changes it to $\frac{n}{m}\in\mathbb{Q}$ (rotation by ±$\pm\frac{\pi}{2}$), both of which are rational; a contradiction.
$\square$
Corollary 3.8
@@
Indeed, both billiard trajectories unfold to the same line with rational slope in the plane $\mathbb{R}^{2}$. Hence, every singular orbit is contained in a billiard trajectory that both starts and terminates at a vertex. Such billiard trajectories are called generalised diagonals [23].
Definition 3.9
@@
Proof. The unfolded trajectory of such a generalised diagonal is the line through the origin with slope $\frac{m}{n}$ that passes through all vertex points $(i,j)\in\mathbb{Z}\times\mathbb{Z}$ that are integer multiples of $(n,m)$. There are no other points $(i,j)\in\mathbb{Z}\times\mathbb{Z}$ on this line, because $\gcd{(m,n)}=1$. From $(0,0)$ to $(n,m)$, we encounter exactly $m-1$ vertical reflections and exactly $n-1$ horizontal reflections, which means that there are $m-1$ collision points on the horizontal sides AB or CD and $n-1$ collision points on the vertical sides BC or DA. The length of this singular trajectory is, thus, $m-1+n-1=m+n-2$, as required.
$\square$
@@
Proof. Since $\tan{(\alpha_{0})}=\frac{p}{q}$ is rational, this billiard trajectory cannot be a non-periodic orbit, because non-periodic orbits unfold to lines with irrational slopes; cf. Remark 2. If it is periodic, then it is of type $(p,q)$ and Definitions 2.2 and 3.4 imply that such a periodic orbit will be a member of the class $\mathcal{C}_{K}(p)$, with $K=p+q$. Furthermore, the billiard trajectory unfolds to a line (or line segment) with fixed slope $\frac{p}{q}$, so if it is singular, it must lie on a generalised diagonal with this slope. Since $\gcd{(p,q)}=2$, Proposition 3.10 implies that the generalised diagonal will have length $\frac{p}{2}+\frac{q}{2}-2$ as required, and it will, indeed, have $\frac{p}{2}-1$ collision points on the horizontal sides AB or CD and the remaining $\frac{q}{2}-1$ collision points on the vertical sides BC or DA. Hence, the proof is complete as soon as we show that a billiard trajectory generated by the pair $\langle P_{0},\alpha_{0}\rangle$ is singular if and only if $P_{0}=\frac{2\ell}{p}$ for $\ell=0,1,\ldots,\frac{p}{2}-1$.
@@ 5 Discussion
We completely classified the existence and nature of all periodic orbits for a rectangular billiard table with aspect ratio $1:\varrho$, where $0<\varrho<\infty$; we discussed the square billiard table in full detail, which is included in this setting as the case with $\varrho=1$. The class of rectangular billiars is special, because it is the only class of polygonal billards for which the classification of all periodic orbits is preserved under a non-uniform scaling [9, 15]. The special property that the sides of a rectangular table are perpendicular to each other enables us to characterise the periodic orbits for this class of billard tables in unprecedented detail. In particular, we show that rectangular billiards admit period-$K$ orbits for any even period, but not for any odd period $K\in\mathbb{N}$. As soon as $K\geq 6$, there are different types of period-$K$ orbits, determined by the difference between the numbers of collisions with the two pairs of parallel sides. Each type is completely determined by the slope that the periodic orbit should have with respect to one of the sides of the rectangle. Furthermore, the periodic orbit can be realised with any initial point on this side, except for a finite set of exactly $\frac{1}{2}K-1$ points, other than the table corners, that give rise to singular orbits; the unfolding of such singular orbits leads to generalised diagonals in the plane. We defined equivalence classes for each type and proved that the total of different types of period-$K$ orbits, for any $K\in 2\,\mathbb{N}$ is given by Euler’s totient function $\varphi(N)$ evaluated at $N=\frac{1}{2}K$; the totients $n\in\{1,\ldots,N\}$ such that $\gcd{(n,N)}=1$ define pairs $(m,n)$ with $m=N-n$ that uniquely define the slopes needed to generate a periodic orbit of a specific type.
@@
As mentioned, the difference between the square and rectangular billiard is a stretching of one pair of parallel sides, which leads to an adjustment by the aspect-ratio parameter $\varrho$ of the slopes required to generate particular periodic orbits. It is tempting to apply a more general transformation, for example, one that scales and shears the sides, such that the billiard table becomes a parallelogram. Unfortunately, the unfolding technique generally fails to tile the plane for parallelogram billiards [16, 41, 39]. We find that there is very little known regarding parallelogram billiards. Do parallelograms also only admit periodic orbits with even periods? What are necessary and sufficient conditions for existence of a period-$K$ orbit and how do we construct such a periodic orbit for the parallelogram?
@@
The non-existence of periodic orbits with odd periods has been proven for parallelograms consisting of two equilateral triangles glued together, that is, one angle is $\frac{\pi}{3}$ radians [2]. Furthermore, it is known that periodic orbits for the parallelogram with angles $\frac{\pi}{4}$ and $\frac{3\pi}{4}$ do not persist under perturbations, that is, so-called stable periodic orbits do not exist for this parallelogram table [32]. More generally, it is known that polygonal billiard tables with angles equal to rational multiples of $\pi$ have many periodic trajectories [7, 12, 28, 39]. However, almost nothing is known for billiard tables with irrational angles. For example, as we already alluded to in the introduction, it is an open question whether periodic orbits exist for triangular billiard tables with arbitrary (irrational) angles. The (rational) angles for acute triangles that lead to a lattice polygon, that is, triangles that unfold to a planar tiling, have all been listed [26, 31]. Hooper [20] has shown that periodic orbits for right-triangle billiards are never stable. In contrast, if one perturbs the right triangle such that it has an obtuse angle, then it will always have a stable periodic orbit [37, 38]; the proof is restricted to obtuse angles that do not exceed $\frac{5\pi}{9}$. For larger angles, we are only aware of existence results of periodic orbits for isosceles triangles [21]. We believe that progress can be made from the study of periodic orbits in a parallelogram table, which we view as continuous deformations from corresponding periodic orbits in the rectangular table. Results in this direction are left for future work.